OpenStudy (anonymous):

PLEASE HELP ME!!!!! The perimeter of a rectangle is 68 ft. The ratio of its length to its width is 9 : 8. What are the dimensions of the rectangle? A. 24 ft by 10 ft B. 20 ft by 14 ft C. 22 ft by 12 ft D. 18 ft by 16 ft

OpenStudy (anonymous):

C

OpenStudy (anonymous):

You need an explanation to why it is C.

OpenStudy (anonymous):

let l be 9x and w be 8x for easy referrence

OpenStudy (anonymous):

Thank you!!

OpenStudy (anonymous):

The only one that maintains the ratio of 9:8. Twice 9 is 18. Twice 8 is 16. Get it?

OpenStudy (anonymous):

its D.

OpenStudy (anonymous):

yeah thanks a million

OpenStudy (anonymous):

No problem.

OpenStudy (anonymous):

can you help with one more

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

Sure.

OpenStudy (anonymous):

What is the ratio of x to y? 3x = 8y A. 1/3 B. 1/8 C. 8/3 D. 3/8

OpenStudy (anonymous):

the answer is C

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

NY MORE QUESTION?

OpenStudy (anonymous):

I believe that "find the ratio of x to y" literally means solving for (x/y). Starting with FutureMathProfessor's result, x = 8/3y, simply divide both sides of the equation by y: x=(8/3)y becomes xy=83.

OpenStudy (anonymous):

no thats it